# One-Shot Imitation under Mismatched Execution

### v5

@@ Abstract
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. However, translating these demonstrations into robot-executable actions presents significant challenges due to execution mismatches in movement styles and physical capabilities. Existing methods either depend on human-robot paired data, which is infeasible to scale, or rely heavily on frame-level visual similarities that often break down in practice. To address these challenges, we propose RHyME, a novel framework that automatically aligns human and robot task executions using optimal transport costs. Given long-horizon robot demonstrations, RHyME synthesizes semantically equivalent human videos by retrieving and composing short-horizon human clips. This approach facilitates effective policy training without the need for paired data. RHyME successfully imitates a range of cross-embodiment demonstrators, both in simulation and with a real human hand, achieving over 50% increase in task success compared to previous methods. We release our code and datasets at this website.
Figure 1: Overview of RHyME. We introduce RHyME, a hierarchical framework that trains a robot policy to mimic a long-horizon video from a demonstrator that exhibits mismatched task execution. Inference Time (Left): Our robot policy translates a demonstrator video into actions to complete the same long-horizon tasks specified by the input. Train Time (Right): Given unpaired robot-demonstrator datasets, RHyME ”imagines” a paired dataset by employing sequence-level similarity metrics which can be used for training the policy.
@@ I Introduction
We propose RHyME (Retrieval for Hybrid Imitation under Mismatched Execution), a framework which trains a robot policy to follow a long-horizon demonstration from a mismatched expert without access to paired human-robot videos (Fig. 1). First, RHyME defines a sequence-level similarity metric between human and robot embeddings, using optimal transport to measure alignment. Given a robot trajectory and a database of human play data, RHyME imagines a long-horizon video by retrieving and composing short-horizon snippets of human demonstrations similar to the robot video. This retrieval process is guided by an optimal transport similarity metric between human and robot sequences. The framework trains a policy using a hybrid approach, incorporating both real robot demonstrations and imagined human sequences. Our contributions can be summarized as:
@@
    We propose RHyME, a retrieval-based algorithm for one-shot imitation of long-horizon demonstrations of humans with mismatched execution.
@@
    Without access to paired human-robot datasets, RHyME uses optimal transport to imagine human demonstrations corresponding to each robot trajectory.
@@
    We release three novel datasets for cross-embodiment learning in simulation with increasing levels of execution mismatch compared to the robot. RHyME outperforms baselines in simulation and the real world, improving task completion by $2$x when imitating unseen human videos.
@@ II Related Work
One-shot Visual Imitation from Demonstration Videos. We tackle this problem setting in our work where the robot imitates actions from human demonstration videos in a one-shot setting, i.e., the robot uses a prompt video as a guide, aiming to replicate the demonstrated actions after viewing them once [7, 8, 4, 6, 41, 42, 43, 32]. If a paired dataset of human and robot videos executing the same task exists, the robot can learn to translate a prompt video into actions directly [7, 8]. The closest to our work is the setting without paired data of human and robot skills. Prior works [6, 4] train policies conditioning on robot videos and zero-shot transfer to a prompt demonstration at test time using aligned visual embeddings. For example, XSkill [4] uses a self-supervised clustering algorithm based on visual similarity to align representations of human and robot videos. However, such an approach can falter when there are significant mismatches in execution. We address this issue by posing the visual imitation problem as a train-time retrieval problem. During training, we match robot videos to the closest human snippets from an unpaired play dataset to imagine synthetic demonstration videos. Training robot policies conditioned on these synthetic videos enable the robot to translate demonstration videos into robot actions.

### v6

@@ Abstract
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. However, translating these demonstrations into robot-executable actions presents significant challenges due to execution mismatches in movement styles and physical capabilities. Existing methods for human-robot translation either depend on paired data, which is infeasible to scale, or rely heavily on frame-level visual similarities that often break down in practice. To address these challenges, we propose RHyME, a novel framework that automatically pairs human and robot trajectories using sequence-level optimal transport cost functions. Given long-horizon robot demonstrations, RHyME synthesizes semantically equivalent human videos by retrieving and composing short-horizon human clips. This approach facilitates effective policy training without the need for paired data. RHyME successfully imitates a range of cross-embodiment demonstrators, both in simulation and with a real human hand, achieving over 50% increase in task success compared to previous methods. We release our code and datasets at this website.
Figure 1: Overview of RHyME. We introduce RHyME, a hierarchical framework that trains a robot policy to mimic a long-horizon video from a demonstrator that exhibits mismatched task execution. Train Goal (Left): Given unpaired human-robot datasets, RHyME “imagines” employs sequence-level similarity functions to create a paired dataset for training one-shot imitation robot policies. Inference Goal (Right): Our robot policy translates a human video into robot actions to perform the specified long-horizon task.
@@ I Introduction
We propose RHyME (Retrieval for Hybrid Imitation under Mismatched Execution), a framework which trains a robot policy to follow a long-horizon demonstration from a mismatched expert without access to paired human-robot videos (Fig. 1). First, RHyME defines a sequence-level similarity metric between human and robot embeddings, using optimal transport to measure alignment. Given a robot trajectory and a database of human play data, RHyME imagines a long-horizon video by retrieving and composing short-horizon snippets of human demonstrations similar to the robot video. This retrieval process is guided by an optimal transport similarity metric between human and robot sequences. The framework trains a policy using a hybrid approach, incorporating both real robot demonstrations and imagined human sequences. Our contributions are:
@@
    We introduce RHyME, a retrieval-based algorithm to create paired datasets of human and robot trajectories, enabling one-shot imitation under mismatched execution.
@@
    We propose three novel datasets for cross-embodiment learning in the Franka Kitchen simulator with progressively increasing execution mismatch relative to the robot.
@@
    In both simulation and the real world, we demonstrate that RHyME outperforms baselines, increase task completion rates $2$x when imitating unseen human videos. Additionally, we show that sequence-level similarity functions significantly improve retrieval performance compared to frame-level similarity functions.
@@ II Related Work
One-shot Visual Imitation from Demonstration Videos. We tackle this problem setting in our work where the robot imitates actions from human demonstration videos in a one-shot setting, i.e., the robot uses a prompt video as a guide, aiming to replicate the demonstrated actions after viewing them once [7, 8, 4, 6, 41, 42, 43, 32]. If a paired dataset of human and robot videos executing the same task exists, the robot can learn to translate a prompt video into actions directly [7, 8]. The closest to our work is the setting without paired data of human and robot skills. However, these works [6, 4] train policies conditioning on robot videos and rely on zero-shot transfer to a prompt demonstration at test time. For example, XSkill [4] uses a self-supervised clustering algorithm based on visual similarity to align representations of human and robot videos. However, such an approach can falter when there are significant mismatches in execution. We address this issue by posing the visual imitation problem as a train-time retrieval problem. During training, we match robot videos to the closest human snippets from an unpaired play dataset to imagine synthetic demonstration videos. Training robot policies conditioned on these synthetic videos enable the robot to translate demonstration videos into robot actions.